OpenStudy (anonymous):

can you break down the steps for this question m - n/9 when m = 42 and n = 54

OpenStudy (anonymous):

it also said to evaluate the expression

Vocaloid (vocaloid):

n = 54 n/9 = ?

OpenStudy (anonymous):

42/9

Vocaloid (vocaloid):

read carefully... n is 54... n/9 = ?

OpenStudy (anonymous):

54/9

OpenStudy (anonymous):

I already have the answer I just need someone to explain the formula

Vocaloid (vocaloid):

ok, the problem tells us that m = 42 and n = 54 m - n/9 means replace "m" with 42 and replace "n" with 54... 42 - 54/9 = 42 - 6 = ?

OpenStudy (anonymous):

36

Vocaloid (vocaloid):

yup that's it

OpenStudy (anonymous):

Thanks